### v1

# Eugene Kritchevski

Abstract
We introduce an elementary framework that captures the mechanism driving the alignment of rows and columns in products of positive matrices. All worst-case misalignment occurs already in dimension two, leading to an explicit collapse principle and a sharp nonlinear bound for finite products. The proof avoids Hilbert-metric and cone-theoretic techniques, relying instead on basic calculus. In the Hilbert metric, the classical Birkhoff–Bushell contraction captures only the linearized asymptotic regime, whereas our nonlinear envelope function gives the exact worst-case behavior for finite products.
1 Introduction
Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all rows and columns become nearly proportional. Classical methods approach this behavior through the Hilbert projective metric, probabilistic tools, or spectral theory. In this work we take a different route: we work directly with the distortion
$R(A):=\max_{i,j,k,\ell}\frac{a_{ik}a_{j\ell}}{a_{i\ell}a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}},$
which quantifies how far the rows or columns of $A$ are from being proportional. We have $R(A)=1$ in the perfectly aligned case; in general $R(A)\geq 1$ and increases with misalignment. Although $R(A)$ corresponds to the exponential of the projective diameter of $A$, we rely only on its elementary ratio form and use no projective-metric theory.
Tracking how distortion changes under matrix multiplication gives an elementary, coordinate-level mechanism for alignment. The central observation is that worst–case distortion occurs already in dimension two; higher dimensions do not produce any new extremal behaviour. In dimension two, the explicit sharp distortion bound can then be computed by an elementary one–variable calculus optimization. A dimension collapse argument shows that exactly the same bound holds in arbitrary dimension. This yields the envelope inequality
$R(AB)\leq\Phi(R(A),R(B)),$
where
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{\!2}.$
Roadmap.
Section 2 introduces distortion and proves the envelope inequality in dimension 2. Section 3 extends the envelope inequality to arbitrary dimension via a collapse argument. Section 4 develops distortion for longer products of matrices, including explicit Frobenius–Perron bounds. Section 5 compares our result with the classical Birkhoff–Bushell theory. Section 6 highlights several directions for further work. Additional calculus lemmas used in the proofs are collected in the Appendix.
2 Distortion Calculus in Dimension 22
Definition 2.1.
Let
$A=\begin{pmatrix}a_{11}&a_{12}\\[2.0pt] a_{21}&a_{22}\end{pmatrix}$
be a $2\times 2$ matrix with strictly positive entries.
  1. 1.
    The oriented distortion of $A$ is
    $F(A):=\frac{a_{11}a_{22}}{\,a_{12}a_{21}\,}.$
  2. 2.
    The distortion of $A$ is
    $R(A):=\max\{F(A),\,1/F(A)\}.$
Both $F(A)$ and $R(A)$ measure how far the two columns (equivalently, the two rows) of $A$ are from being proportional. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate from the definition:
  • Scaling invariance. Multiplying any row or any column of $A$ by a positive scalar leaves both $F(A)$ and $R(A)$ unchanged.
  • Effect of swapping. Swapping the two rows or the two columns of $A$ replaces $F(A)$ by $1/F(A)$ and therefore leaves $R(A)$ unchanged.
  • Orientation via determinant. For $2\times 2$ positive matrices,
    $R(A)=F(A)\iff F(A)\geq 1\iff\det(A)\geq 0,$
    and
    $R(A)=1/F(A)\iff F(A)\leq 1\iff\det(A)\leq 0.$
    In the perfectly aligned case, $\det(A)=0$ and both orientations coincide. Note that in our projective, scale-invariant setting, the area interpretation of the determinant is irrelevant; only the sign information matters.
Our goal is to obtain a sharp explicit upper bound on the distortion
$R(AB)$
of the product of two positive $2\times 2$ matrices. The analysis is greatly simplified by a few distortion-preserving normalizations.
We first ensure positive orientation:
$F(A)\geq 1,\qquad F(B)\geq 1,$
so that $F(A)=R(A)$ and $F(B)=R(B)$. This can always be achieved by swapping the rows of $A$ and/or the columns of $B$; such swaps do not affect the distortion of $A$, $B$, or $AB$. After ensuring positive orientation of $A$ and $B$, we have $\det(A)\geq 0$, $\det(B)\geq 0$, and so $\det(AB)=\det(A)\det(B)\geq 0$. Then $F(AB)\geq 1$ and $AB$ is also positively oriented: $F(AB)=R(AB)$.
Next, we are free to multiply the rows of $A$ and the columns of $B$ by any positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
After these normalizations, we may assume, without loss of generality, that $A$ and $B$ are in the canonical form
$A=\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix},\qquad B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix}$,
where
$\alpha=R(A)\geq 1,\qquad\beta=R(B)\geq 1.$
A direct computation gives
$AB=\begin{pmatrix}1+uv&1+\beta\,uv\\[5.69054pt] 1+\alpha\,uv&1+\alpha\beta\,uv\end{pmatrix}=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix}$,
where
$t=uv>0.$
Then
$R(AB)=f_{\alpha,\beta}(uv),$
where
$f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)}\qquad t>0.$
Thus the problem of finding a sharp upper bound for $R(AB)$ in terms of $R(A)$ and $R(B)$ reduces to a simple calculus exercise: it suffices to maximize the function $f_{\alpha,\beta}(t)$ for $t>0$. The details are carried out in Lemma A.1 and the result is
$\max_{t>0}f_{\alpha,\beta}(t)=\Phi(\alpha,\beta),$
where
$\Phi(\alpha,\beta)=\frac{\bigl(1+\sqrt{\alpha\beta}\bigr)^{2}}{\bigl(\sqrt{\alpha}+\sqrt{\beta}\bigr)^{2}}.$
We refer to $\Phi$ as the envelope function.
Thus we obtain
$R(AB)=f_{\alpha,\beta}(uv)\leq\Phi(\alpha,\beta),$
which proves the following bound:
Theorem 2.2 (Envelope Inequality in Dimension $2$).
Let $A$ and $B$ be strictly positive $2\times 2$ matrices. Then
$\boxed{R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}}.$
Moreover, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable (Lemma A.2). Therefore, from Theorem 2.2 we obtain:
Corollary 2.3.
Let $A$ and $B$ be strictly positive $2\times 2$ matrices, and suppose
$R(A)\leq\alpha,\qquad R(B)\leq\beta.$
Then
$R(AB)\leq\Phi(\alpha,\beta).$
We note that the envelope inequality is not merely an upper bound; the maximal distortion is attainable.
Proposition 2.4 (Two–Sided Sharpness).
Let $\alpha,\beta\geq 1$.
  1. (i)
    Given any $A>0$ with $R(A)=\alpha$, one can choose $B>0$ with $R(B)=\beta$ such that
    $R(AB)=\Phi(\alpha,\beta).$
  2. (ii)
    Given any $B>0$ with $R(B)=\beta$, one can choose $A>0$ with $R(A)=\alpha$ such that
    $R(AB)=\Phi(\alpha,\beta).$
Proof.
We first prove (i). By Lemma A.1, the function $f_{\alpha,\beta}(t)$ attains its maximum $\Phi(\alpha,\beta)$ at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}}.$
Every product of the form
$\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad uv=t^{*},$
has distortion $\Phi(\alpha,\beta)$.
Let $A>0$ be given with $R(A)=\alpha$. By rescaling the rows of $A$ by positive factors (which does not change $R(A)$ nor $R(AB)$), we may assume that
$A=\begin{pmatrix}1&x\\[5.69054pt] 1&y\end{pmatrix},\qquad x,y>0.$
If $A$ is positively oriented, i.e. $y/x=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
If $A$ is negatively oriented, i.e. $x/y=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
In both cases one checks directly that
$R(AB)=\Phi(\alpha,\beta).$
This proves (i).
For (ii), use that $R(M)=R(M^{\mathsf{T}})$ and $\Phi(\alpha,\beta)=\Phi(\beta,\alpha)$. Given $B$ with $R(B)=\beta$, apply (i) to $B^{\mathsf{T}}$ and $\alpha$: there exists $C>0$ with $R(C)=\alpha$ such that
$R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha).$
Setting $A=C^{\mathsf{T}}$ gives
$R(AB)=R((AB)^{\mathsf{T}})=R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha)=\Phi(\alpha,\beta),$
which proves (ii). ∎
Corollary 2.5 (optimization identity).
For all $\alpha,\beta\geq 1$,
$\boxed{\Phi(\alpha,\beta)=\max\bigl\{\,R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\,\bigr\}}.$
3 Reduction to Two Dimensions in the General Case
The goal of this section is to extend the envelope inequality to arbitrary dimension.
For $x,y\in\mathbb{R}_{>0}^{d}$, define
$s^{+}(x,y)=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y)=\min_{i}\frac{y_{i}}{x_{i}},$
and the distortion
$\operatorname{Dist}(x,y)=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
For perfectly aligned vectors we have $s^{+}(x,y)=s^{-}(x,y)$, so $\operatorname{Dist}(x,y)=1$. When the ratios $y_{i}/x_{i}$ are more scattered, the distortion increases. Figure 1 illustrates $s^{+}(x,y)$ and $s^{-}(x,y)$ as maximal and minimal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
Figure 1: The pair $(x,y)$ represented as a cloud of points between the extremal rays of slopes $s^{+}(x,y)$ and $s^{-}(x,y)$ in $\mathbb{R}_{>0}^{2}$.
We extend the definition of distortion to positive matrices of size $d_{1}\times d_{2}$ by
$R(A)=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
Thus $R(A)$ is the maximum of the distortions of all $2\times 2$ submatrices of $A$, obtained by choosing any two rows and any two columns. Equivalently, it is the maximum distortion between any two columns (or two rows) of $A$:
$R(A)=\max_{k,\ell}\operatorname{Dist}\bigl(A_{*k},A_{*\ell}\bigr)=\max_{i,j}\operatorname{Dist}\bigl(A_{i*},A_{j*}\bigr).$
Define $F_{d}:(\mathbb{R}^{d}_{>0})^{4}\to\mathbb{R}_{>0}$ by
$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)}.$
Here $x\cdot u$ denotes the standard Euclidean dot product.
Lemma 3.1 (coordinatewise extremal ratio pushing).
Let $d\geq 2$ and let $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Define
$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
Then there exist vectors $y^{*},v^{*}\in\mathbb{R}^{d}_{>0}$ such that
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad s^{\pm}(u,v^{*})=s^{\pm}(u,v),$
and for all $i$,
$\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\},\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}.$
Proof.
We will first construct $y^{*}$ and then use a similar procedure to obtain $v^{*}$. The strategy is to adjust the coordinates of $y$ one at a time, always modifying them in a way that does not decrease the value of $F_{d}(x,y;u,v)$, and while keeping the extremal slopes $s^{\pm}(x,y)$ unchanged.
To make this precise, partition the index set $\{1,\dots,d\}$ into
$I_{-}:=\{\,i:y_{i}/x_{i}=a^{-}\,\},\qquad I_{+}:=\{\,i:y_{i}/x_{i}=a^{+}\,\},$
and
$I_{0}:=\{1,\dots,d\}\setminus(I_{-}\cup I_{+}).$
We will never alter the coordinates in $I_{-}$ or $I_{+}$, so that the extremal slopes remain fixed at all times. Thus we immediately set
$y_{i}^{*}:=y_{i}\qquad\text{for all }i\in I_{-}\cup I_{+}.$
The coordinatewise replacement procedure is applied only to indices in $I_{0}$.
We begin with a simple coordinate monotonicity observation. Fix $x,u,v$ and all coordinates of $y$ except $y_{i}$. Then
$f(y_{i}):=F_{d}(x,y;u,v)$
is a fractional linear function of $y_{i}$ with strictly positive denominator, and is therefore monotone (either increasing or decreasing) in $y_{i}$. If $f$ is increasing in $y_{i}$, we set
$y_{i}^{*}:=a^{+}\,x_{i},$
while if $f$ is decreasing in $y_{i}$, we set
$y_{i}^{*}:=a^{-}\,x_{i}.$
Then
$F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),\qquad s^{\pm}(x,y^{*})=s^{\pm}(x,y),$
and
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
The key point is that replacing $(x_{i},y_{i})$ by $(x_{i},y_{i}^{*})$ does not decrease $F_{d}$, does not change $s^{\pm}(x,y)$, and forces the ratio $y_{i}^{*}/x_{i}$ to be extremal. The frozen extremal indices in $I_{-}\cup I_{+}$ ensure that the original extremal ratios are preserved throughout.
We now apply this replacement successively to all $i\in I_{0}$, always leaving previously modified coordinates unchanged. This yields a vector $y^{*}$ such that
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),$
and every ratio $y_{i}^{*}/x_{i}$ is extremal:
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
With $y^{*}$ fixed, we now apply exactly the same coordinatewise extremal-pushing procedure to the pair $(u,v)$. We freeze the indices where the ratios $v_{i}/u_{i}$ already attain the extremal values $b^{\pm}$, and modify only the remaining coordinates exactly as above: for each active coordinate we replace $v_{i}$ by either $b^{-}u_{i}$ or $b^{+}u_{i}$, according to the monotonicity of $F_{d}(x,y^{*};u,v)$ in $v_{i}$. This produces a vector $v^{*}$ such that
$s^{\pm}(u,v^{*})=s^{\pm}(u,v),\qquad F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y^{*};u,v),$
and each ratio $v_{i}^{*}/u_{i}$ is extremal:
$\frac{v_{i}^{*}}{u_{i}}\in\{\,b^{-},\,b^{+}\,\}.$
This completes the proof. ∎
Lemma 3.2 ($2$-point support collapse).
Assume that the coordinate ratios of $y^{*}/x$ take only the two extremal values $a^{\pm}$, and the coordinate ratios of $v^{*}/u$ take only the two extremal values $b^{\pm}$, where
$a^{\pm}=s^{\pm}(x,y^{*}),\qquad b^{\pm}=s^{\pm}(u,v^{*}).$
Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;F_{d}(x,y^{*};u,v^{*}),$
and
$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}.$
Proof.
If $a^{+}=a^{-}=a$, then $y^{*}=ax$ and hence $F_{d}(x,y^{*};u,v^{*})=1$. The lemma then holds, for example, with
$x^{\prime}=(1,1),\qquad y^{\prime}=(a,a),\qquad u^{\prime}=(1,1),\qquad v^{\prime}=(b^{-},b^{+}).$
We now treat the nontrivial case $a^{+}>a^{-}$. Define
$I_{+}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{+}\bigr\},\qquad I_{-}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{-}\bigr\},$
which form a partition of $\{1,\dots,d\}$. Set
$x_{\pm}:=\sum_{i\in I_{\pm}}x_{i},\qquad u_{\pm}:=\frac{1}{x_{\pm}}\sum_{i\in I_{\pm}}x_{i}u_{i},$
so that
$x_{\pm}u_{\pm}=\sum_{i\in I_{\pm}}x_{i}u_{i}.$
Set
$y_{\pm}:=a^{\pm}x_{\pm},\qquad v_{\pm}:=b^{\pm}u_{\pm},$
and define the two-dimensional vectors
$x^{\prime}:=(x_{-},x_{+}),\qquad y^{\prime}:=(y_{-},y_{+}),\qquad u^{\prime}:=(u_{-},u_{+}),\qquad v^{\prime}:=(v_{-},v_{+}).$
A direct computation using the above definitions gives
$x^{\prime}\cdot u^{\prime}=x_{-}u_{-}+x_{+}u_{+}=\sum_{i\in I_{-}}x_{i}u_{i}+\sum_{i\in I_{+}}x_{i}u_{i}=x\cdot u,$
and similarly
$x^{\prime}\cdot v^{\prime}=x\cdot v^{*},\qquad y^{\prime}\cdot u^{\prime}=y^{*}\cdot u,\qquad y^{\prime}\cdot v^{\prime}=y^{*}\cdot v^{*}.$
Hence
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=\frac{(x^{\prime}\cdot u^{\prime})(y^{\prime}\cdot v^{\prime})}{(x^{\prime}\cdot v^{\prime})(y^{\prime}\cdot u^{\prime})}=\frac{(x\cdot u)(y^{*}\cdot v^{*})}{(x\cdot v^{*})(y^{*}\cdot u)}=F_{d}(x,y^{*};u,v^{*}),$
which completes the proof.
Corollary 3.3 (Four-Point Collapse).
Let $d\geq 2$ and $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Then there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$s^{\pm}(x^{\prime},y^{\prime})=s^{\pm}(x,y),\qquad s^{\pm}(u^{\prime},v^{\prime})=s^{\pm}(u,v),$
and
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
Proof.
Apply lemma 3.1 to obtain $y^{*},v^{*}$ with
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
then lemma 3.2 to $(x,y^{*};u,v^{*})$ to obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=F_{d}(x,y^{*};u,v^{*}),$
which gives the claim. ∎
Theorem 3.4 (Envelope Inequality in Arbitrary Dimension).
Let $A\in\mathbb{R}^{d_{1}\times d_{2}}_{>0}$ and $B\in\mathbb{R}^{d_{2}\times d_{3}}_{>0}$, where $d_{1},d_{2},d_{3}\geq 2$. Then
$\boxed{\,R(AB)\;\leq\;\Phi\big(R(A),R(B)\big)\,}.$
Proof.
By definition, $R(AB)$ is the maximum distortion among all $2\times 2$ sub-blocks of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ such that this maximum is attained.
Let
[…]
Then the corresponding $2\times 2$ block of $AB$ is
$\begin{pmatrix}x\cdot u&x\cdot v\\ y\cdot u&y\cdot v\end{pmatrix}$,
whose distortion is
$R(AB)\;=\;F_{d}(x,y;u,v).$
Set
$\alpha:=\operatorname{Dist}(x,y)\leq R(A),\qquad\beta:=\operatorname{Dist}(u,v)\leq R(B).$
By Corollary 3.3, there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
$\operatorname{Dist}(x^{\prime},y^{\prime})=\operatorname{Dist}(x,y)=\alpha,\qquad\operatorname{Dist}(u^{\prime},v^{\prime})=\operatorname{Dist}(u,v)=\beta,$
and
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then $R(A^{\prime})=\operatorname{Dist}(x,y)=\alpha$ and $R(B^{\prime})=\operatorname{Dist}(u,v)=\beta$. Then
$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix}$,
and therefore
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;R(A^{\prime}B^{\prime}).$
Combining the above steps and using that $\Phi$ is increasing in each variable gives
$R(AB)$ $=F_{d}(x,y;u,v)$
$\leq F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$
$=R(A^{\prime}B^{\prime})$
$\leq\Phi(\alpha,\beta)$
$\leq\Phi(R(A),R(B)).$
This completes the proof.
4 Distortion of Products
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and define the partial products
$P_{n}:=A_{n}\cdots A_{1}.$
The matrices may have arbitrary dimensions, provided the products are well defined.
It is convenient to work with the square root of the distortion,
$S(A):=\sqrt{R(A)}.$
The envelope inequality then takes the simpler form
$S(AB)\;\leq\;\Psi\bigl(S(A),S(B)\bigr),\qquad\Psi(p,q):=\frac{1+pq}{p+q}.$
Since $\Psi$ is increasing in each variable, distortion bounds propagate along the product. If
$R(A_{n})\leq\alpha_{n},\qquad\text{i.e.}\qquad S(A_{n})\leq p_{n}:=\sqrt{\alpha_{n}},$
then
$S(P_{n})\leq q_{n},$
where $(q_{n})$ is defined recursively by
$q_{1}=p_{1},\qquad q_{n+1}=\Psi(p_{n+1},q_{n}).$
Finally,
$R(P_{n})\leq q_{n}^{2}.$
Uniformly bounded distortions.
Assume that all matrices satisfy the uniform bound
$R(A_{n})\leq\alpha,\qquad\text{i.e.}\qquad S(A_{n})\leq p:=\sqrt{\alpha}.$
Then the upper bounds $q_{n}$ satisfy
$q_{n+1}=\Psi(p,q_{n}),\qquad q_{1}=p.$
With $p$ fixed, this iterates a single Möbius map. Writing
$\kappa:=\frac{p-1}{p+1},$
its $n$th iterate can be expressed as
$q_{n}=1+\frac{2\,\kappa^{n}}{1-\kappa^{n}}.$
Consequently,
$R(P_{n})\leq q_{n}^{2}=1+\frac{4\,\kappa^{n}}{(1-\kappa^{n})^{2}}.$
For convenience, we record the following standalone theorem.
Theorem 4.1 (Finite-Product Distortion Bound).
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices, possibly of varying dimensions, and assume that all products $P_{n}:=A_{n}\cdots A_{1}$ are well defined.
For any positive matrix $B>0$, define its distortion by
$R(B):=\max_{i,j,k,\ell}\frac{b_{ik}\,b_{j\ell}}{b_{i\ell}\,b_{jk}}.$
Assume that all factors satisfy the uniform bound
$R(A_{n})\leq\alpha\qquad\text{for all }n\geq 1.$
Set
$\kappa:=\frac{\sqrt{\alpha}-1}{\sqrt{\alpha}+1}.$
Then, for every $n\geq 1$,
$R(P_{n})\;\leq\;1\;+\;\frac{4\,\kappa^{\,n}}{(1-\kappa^{\,n})^{2}},$
and therefore
$R(P_{n})\longrightarrow 1\quad\text{at least exponentially fast as }n\to\infty.$
5 Comparison with the Birkhoff–Bushell
Contraction Theory
In classical Birkhoff–Bushell (BB) theory one studies distortion on a log scale via the Hilbert metric
$d_{H}(x,y)=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
If $B$ is an $n\times 2$ matrix with columns $x$ and $y$, and $A$ is any positive $n\times n$ matrix, then
$d_{H}(x,y)=\log R(B),\qquad d_{H}(Ax,Ay)=\log R(AB).$
In this log setting, our envelope inequality becomes
$d_{H}(Ax,Ay)\;\leq\;\Theta\!\left(d_{H}(x,y)\right),$
where
$\Theta(h)=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p=\sqrt{R(A)}.$
In the nontrivial case $p>1$, a direct computation shows that $\Theta$ is strictly increasing, strictly concave, and satisfies
$\Theta(0)=0,\qquad\Theta^{\prime}(0)=\frac{p-1}{p+1}=:\kappa(A).$
Thus $\Theta$ lies below its tangent line at the origin:
$\Theta(h)\;\leq\;\kappa(A)\,h,$
which yields the classical BB contraction:
$d_{H}(Ax,Ay)\;\leq\;\kappa(A)\,d_{H}(x,y).$
Note that
$\lim_{h\to\infty}\Theta(h)=\log R(A),$
so the envelope bound $\Theta(h)$ remains uniformly bounded, whereas the BB linear bound
$\kappa(A)\,h$
diverges. Thus both bounds agree to first order at $h=0$ (and therefore yield the same Lyapunov–gap asymptotics for distortion of high powers of $A$), but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp in the non-asymptotic regime.
For illustration, Figure 2 compares the nonlinear envelope map $h\mapsto\Theta(h)$ with the classical BB line $k(A)h$, together with the saturation level $\log R(A)$.
[…]
Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$. The envelope contraction saturates at $\log R(A)$, while the BB line grows without bound, showing that the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$ for large $h$.
Historical perspective.
Birkhoff’s fundamental insight was that positive operators act as contractions in the Hilbert metric, allowing the Frobenius–Perron alignment phenomenon to be deduced from the Banach fixed-point theorem. Bushell later refined this viewpoint by computing an explicit contraction constant. In most contraction theorems such constants are technically useful but algebraically unremarkable: they serve as quantitative bounds and carry little intrinsic meaning. Bushell’s constant is a striking exception. Its closed form,
$\kappa(A)=\frac{\sqrt{R(A)}-1}{\sqrt{R(A)}+1},$
is unexpectedly symmetric and rigid—far too elegant, in retrospect, to be merely a sharp Lipschitz coefficient. Moreover, the metric framework of the Hilbert metric plays no essential role in the classical proof: it is used only to invoke the Banach fixed-point theorem, not to explain the mechanism of alignment itself. This indicates that the contraction phenomenon behind the Frobenius–Perron theory is not inherently metric but fundamentally algebraic, encoded in the behaviour of ratios rather than in the geometry induced by the Hilbert metric.
This motivates our approach: the envelope inequality shows that worst-case distortion evolves under Möbius iterations with a common attracting fixed point at $1$. In this formulation, alignment arises from the explicit fixed-point dynamics of these maps rather than from a geometric contraction in the Hilbert metric.
6 Discussion and outlook
The envelope inequality is dimension-free, and the same collapse mechanism extends to a broad class of infinite matrices and positive operators on cones. The operator-theoretic generalizations will be developed elsewhere.
A second direction concerns other projective quantities. Since the collapse argument in Section 3 uses only the four-point ratios
$\frac{(x\!\cdot\!u)(y\!\cdot\!v)}{(x\!\cdot\!v)(y\!\cdot\!u)},$
it is natural to conjecture that analogous two-dimensional collapse mechanisms may apply to a broader class of projective functionals.
The envelope inequality also suggests applications in Markov chain theory. Classical asymptotic mixing bounds typically rely on Doeblin, Dobrushin, or Hilbert-metric methods, whereas the envelope provides an exact non-asymptotic contraction profile. This may offer an alternative route to quantitative bounds for finite-time mixing of positive stochastic matrices.
The recursion in Section 4 provides a deterministic model for worst-case distortion in products of positive matrices; its transfer-matrix form makes the evolution completely explicit. It may also serve as a foundation for combining our methods with probabilistic techniques to study non-asymptotic behavior of random products.
These directions lie beyond the scope of the present paper. Our purpose here is purely foundational: to isolate and prove the core envelope inequality in its simplest form.
Appendix A Calculus Tools
In this appendix we prove properties of the functions (for $\alpha,\beta\geq 1$)
$f_{\alpha,\beta}(t)=\frac{(1+t)(1+\alpha\beta\,t)}{(1+\alpha t)(1+\beta t)},\qquad t>0,$
and
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
Lemma A.1.
The function $f_{\alpha,\beta}$ has the following properties:
  1. 1.
    $f_{\alpha,\beta}(t)\geq 1$ for all $t>0$;
  2. 2.
    if $\alpha,\beta>1$, the unique global maximum is attained at
    $t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
    and
    $f_{\alpha,\beta}(t)\leq f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta);$
  3. 3.
    if $\alpha=1$ or $\beta=1$, then $f_{\alpha,\beta}(t)\equiv 1$.
Proof.
(1) Expanding $f_{\alpha,\beta}(t)\geq 1$ gives
$(1+t)(1+\alpha\beta t)\;\geq\;(1+\alpha t)(1+\beta t),$
which simplifies to
$(\alpha-1)(\beta-1)t\geq 0,$
valid for all $\alpha,\beta\geq 1$ and $t>0$.
(2) Set $\varphi(t)=\log f_{\alpha,\beta}(t)$. A straightforward algebraic simplification yields
$\varphi^{\prime}(t)=\,\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
Thus $\varphi^{\prime}(t)=0$ exactly at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
and this is the unique global maximum. Substituting $t^{*}$ gives the stated value of $f_{\alpha,\beta}(t^{*})$.
(3) If $\alpha=1$ or $\beta=1$, the formula reduces to $f_{\alpha,\beta}(t)\equiv 1$. ∎
Lemma A.2.
For $\alpha,\beta\geq 1$, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable: for all $\alpha_{1}\leq\alpha_{2}$ and $\beta_{1}\leq\beta_{2}$,
$\Phi(\alpha_{1},\beta)\leq\Phi(\alpha_{2},\beta),\qquad\Phi(\alpha,\beta_{1})\leq\Phi(\alpha,\beta_{2}).$
Proof.
Write $p=\sqrt{\alpha}$ and $q=\sqrt{\beta}$, so $p,q\geq 1$, and define
$\psi(p,q)=\frac{1+pq}{p+q},$
so that $\Phi(\alpha,\beta)=\psi(p,q)^{2}$. For fixed $q\geq 1$ we compute
$\frac{\partial\psi}{\partial p}=\frac{q(p+q)-(1+pq)}{(p+q)^{2}}=\frac{q^{2}-1}{(p+q)^{2}}\;\geq\;0.$
Thus $\psi$ is increasing in $p$, and therefore so is $\Phi=\psi^{2}$. By symmetry the same argument applies to $q$. ∎