### v2

# Eugene Kritchevski

Abstract
We study the deviation from proportionality of rows and columns in products of positive matrices. We prove a sharp, dimension-free bound showing that worst-case misalignment is already captured in dimension two and follows an explicit Möbius law, refining the classical Birkhoff–Bushell contraction theory.
1 Introduction
Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all columns (respectively, all rows) become nearly proportional. While alignment is well understood asymptotically, the finite-time law governing worst-case misalignment has not yet been isolated.
In this paper, we develop an elementary framework based on the distortion functional
$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}}.$
Our main result is a sharp envelope inequality controlling the distortion of matrix products,
$R(AB)\leq\Phi(R(A),R(B)),\qquad\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
A key structural observation is a rigidity phenomenon for projective distortion: extremal misalignment does not engage genuinely high-dimensional geometry, but collapses to a two-dimensional extremal configuration.
While distortion is classically measured by the projective diameter $\log R(A)$ in the Hilbert metric, we show that $\sqrt{R(A)}$ is the natural scale on which extremal distortion evolves according to an explicit Möbius law.
Organization of the paper.
Section 2 establishes the envelope inequality in dimension 2. Section 3 extends the result to arbitrary dimension via a dimension-reduction argument. Section 4 derives explicit finite-time bounds for the distortion of matrix products. Section 5 compares the envelope bound with the classical Birkhoff–Bushell theory.
2 Distortion Bounds in Dimension 22
Definition 2.1.
Let
$A=\begin{pmatrix}a_{11}&a_{12}\\[2.0pt] a_{21}&a_{22}\end{pmatrix}$
be a $2\times 2$ matrix with strictly positive entries.
  1. 1.
    The oriented distortion of $A$ is
    $F(A):=\frac{a_{11}a_{22}}{a_{12}a_{21}}.$
  2. 2.
    The distortion of $A$ is
    $R(A):=\max\{F(A),\,1/F(A)\}.$
Both $F(A)$ and $R(A)$ measure the deviation of the two columns (or rows) of $A$ from proportionality. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate.
  • Scaling invariance. Multiplying any row or column of $A$ by a positive scalar leaves $F(A)$ and $R(A)$ unchanged.
  • Effect of swapping. Swapping the two rows or the two columns replaces $F(A)$ by $1/F(A)$ and leaves $R(A)$ unchanged.
  • Orientation via determinant. For $2\times 2$ positive matrices,
    $R(A)=F(A)\iff F(A)\geq 1\iff\det(A)\geq 0,$
    and
    $R(A)=1/F(A)\iff F(A)\leq 1\iff\det(A)\leq 0.$
    In the proportional case $\det(A)=0$, both orientations coincide.
Our goal is to obtain a sharp explicit upper bound on $R(AB)$ for the product of two positive $2\times 2$ matrices. The analysis is simplified by distortion-preserving normalizations.
We first ensure positive orientation,
$F(A)\geq 1,\qquad F(B)\geq 1,$
so that $F(A)=R(A)$ and $F(B)=R(B)$. This can be achieved by swapping rows of $A$ and/or columns of $B$, which does not affect the distortions of $A$, $B$, or $AB$. Then $\det(A),\det(B)\geq 0$, and hence $\det(AB)\geq 0$, so that $F(AB)=R(AB)$.
We may also multiply the rows of $A$ and the columns of $B$ by arbitrary positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
After these normalizations, we may assume without loss of generality that
$A=\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix},\qquad B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix}$,
where
$\alpha=R(A)\geq 1,\qquad\beta=R(B)\geq 1.$
Then
$AB=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix},\qquad t=uv>0.$
Accordingly,
$R(AB)=f_{\alpha,\beta}(t),\qquad f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)},\quad t>0.$
The task of bounding $R(AB)$ therefore reduces to maximizing the one-variable function $f_{\alpha,\beta}(t)$. Differentiating its logarithm yields
$\frac{d}{dt}\log f_{\alpha,\beta}(t)=\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
Thus $f_{\alpha,\beta}$ is maximized at
$t^{*}:=\frac{1}{\sqrt{\alpha\beta}},$
with
$f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta),$
where
$\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
We refer to $\Phi(\alpha,\beta)$ as the envelope function. Hence
$R(AB)\leq\Phi(\alpha,\beta),$
and we obtain the following result.
Theorem 2.2 (Envelope Inequality in Dimension $2$).
Let $A$ and $B$ be strictly positive $2\times 2$ matrices. Then
$R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}.$
The envelope function $\Phi$ is increasing in each variable, yielding the following corollary.
Corollary 2.3.
Let $A$ and $B$ be strictly positive $2\times 2$ matrices with $R(A)\leq\alpha$ and $R(B)\leq\beta$. Then
$R(AB)\leq\Phi(\alpha,\beta).$
The envelope inequality is sharp.
Proposition 2.4 (Two–Sided Sharpness).
Let $\alpha,\beta\geq 1$.
  1. (i)
    For any $A>0$ with $R(A)=\alpha$, there exists $B>0$ with $R(B)=\beta$ such that $R(AB)=\Phi(\alpha,\beta)$.
  2. (ii)
    For any $B>0$ with $R(B)=\beta$, there exists $A>0$ with $R(A)=\alpha$ such that $R(AB)=\Phi(\alpha,\beta)$.
Proof.
We first prove (i). Every product of the form
$\begin{pmatrix}1&u\\ 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad uv=t^{*},$
has distortion $\Phi(\alpha,\beta)$.
Let $A>0$ satisfy $R(A)=\alpha$. By rescaling rows, we may write
$A=\begin{pmatrix}1&x\\ 1&y\end{pmatrix},\qquad x,y>0.$
If $A$ is positively oriented, so that $y/x=\alpha$, we take
$B=\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
If $A$ is negatively oriented, so that $x/y=\alpha$, we take
$B=\begin{pmatrix}1&1\\ \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
In both cases, $R(AB)=\Phi(\alpha,\beta)$, proving (i).
For (ii), we use the symmetry under transposition. Note that
$R(M)=R(M^{\mathsf{T}})\qquad\text{and}\qquad\Phi(\alpha,\beta)=\Phi(\beta,\alpha).$
Given $B>0$ with $R(B)=\beta$, apply part (i) to $B^{\mathsf{T}}$. There exists $C>0$ with $R(C)=\alpha$ such that
$R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha).$
Setting $A=C^{\mathsf{T}}$ gives
$R(AB)=R((AB)^{\mathsf{T}})=R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha)=\Phi(\alpha,\beta),$
which proves (ii). ∎
Corollary 2.5 (Optimization Identity).
For all $\alpha,\beta\geq 1$,
$\Phi(\alpha,\beta)=\max\bigl\{R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\bigr\}.$
3 Reduction to Two Dimensions
In this section we extend the notion of distortion to arbitrary dimension and prove the envelope inequality via a reduction to two dimensions.
For $x,y\in\mathbb{R}_{>0}^{d}$, define
$s^{+}(x,y):=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y):=\min_{i}\frac{y_{i}}{x_{i}}.$
Figure 1 illustrates $s^{\pm}(x,y)$ as the extremal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
Figure 1: Geometric representation of $(x,y)$ as points in $\mathbb{R}_{>0}^{2}$.
The misalignment of $x$ and $y$ is measured by the distortion
$\operatorname{Dist}(x,y):=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
Equivalently,
$\operatorname{Dist}(x,y)=\max_{i,j}\frac{y_{i}x_{j}}{y_{j}x_{i}}.$
We extend the definition of distortion to positive matrices of size $d_{1}\times d_{2}$ by
$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
Thus $R(A)$ is the maximum distortion of all $2\times 2$ submatrices of $A$. Equivalently,
$R(A)=\max_{k,\ell}\operatorname{Dist}\bigl(A_{*k},A_{*\ell}\bigr)=\max_{i,j}\operatorname{Dist}\bigl(A_{i*},A_{j*}\bigr).$
The following functional encodes the oriented distortion of the $2\times 2$ block of a matrix product, obtained by pairing rows $x,y$ of the left factor with columns $u,v$ of the right factor.
Define $F_{d}:(\mathbb{R}^{d}_{>0})^{4}\to\mathbb{R}_{>0}$ by
$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)},$
where $\cdot$ denotes the standard Euclidean dot product.
Lemma 3.1 (coordinatewise extremal ratio pushing).
Let $x,y,u,v\in\mathbb{R}^{d}_{>0}$ and set
$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
Then there exist vectors $y^{*},v^{*}\in\mathbb{R}^{d}_{>0}$ such that
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad s^{\pm}(u,v^{*})=s^{\pm}(u,v),$
and for all $1\leq i\leq d$,
$\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\},\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}.$
Proof.
We first construct $y^{*}$. The strategy is to use monotonicity to adjust the coordinates of $y$ one at a time without decreasing the value of $F_{d}$.
Define
$f(z_{1},\dots,z_{d})\;:=\;F_{d}(x,z;u,v),\qquad z=(z_{1},\dots,z_{d})\in\mathbb{R}_{>0}^{d}.$
For $i\in\{1,\dots,d\}$, write $z_{\neq i}:=(z_{1},\dots,z_{i-1},z_{i+1},\dots,z_{d})$ and
$f_{i}(t\mid z_{\neq i})\;:=\;f(z_{1},\dots,z_{i-1},t,z_{i+1},\dots,z_{d}),\qquad t>0.$
For fixed $z_{\neq i}$, the function $t\mapsto f_{i}(t\mid z_{\neq i})$ is fractional linear with strictly positive denominator, and hence is either increasing, decreasing, or constant.
Set
$B:=\{\,z\in\mathbb{R}_{>0}^{d}:a^{-}x_{i}\leq z_{i}\leq a^{+}x_{i}\text{ for all }i\,\}.$
For each $i\in\{1,\dots,d\}$, define the coordinate update map $E_{i}$ on $B$ by
[…]
Then
$E_{i}(z)\in\{a^{-}x_{i},a^{+}x_{i}\}\quad\text{for all }z\in B.$
Define the associated vector update operator $\mathcal{E}_{i}:B\to B$ by
$(\mathcal{E}_{i}z)_{j}:=\begin{cases}z_{j},&j\neq i,\\ E_{i}(z),&j=i.\end{cases}$
By construction,
[…]
Define
$y^{*}\;:=\;(\mathcal{E}_{d}\circ\cdots\circ\mathcal{E}_{1})(y).$
Then
$f(y^{*})\;\geq\;f(y),\qquad s^{\pm}(x,y^{*})\;=\;s^{\pm}(x,y),\qquad\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\}\ \ \text{for all }i.$
We next construct $v^{*}$ in the same manner. Define
$g(w_{1},\dots,w_{d})\;:=\;F_{d}(x,y^{*};u,w),\qquad w=(w_{1},\dots,w_{d})\in\mathbb{R}_{>0}^{d},$
and
$g_{i}(t\mid w_{\neq i})\;:=\;g(w_{1},\dots,w_{i-1},t,w_{i+1},\dots,w_{d}),\qquad t>0.$
For fixed $w_{\neq i}$, the function $t\mapsto g_{i}(t\mid w_{\neq i})$ is fractional linear with strictly positive denominator. Repeating the coordinate update argument in this setting produces $v^{*}\in\mathbb{R}_{>0}^{d}$ such that
$g(v^{*})\;\geq\;g(v),\qquad s^{\pm}(u,v^{*})\;=\;s^{\pm}(u,v),\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}\ \ \text{for all }i.$
Finally, by construction
$F_{d}(x,y^{*};u,v^{*})$ $=g(v^{*})$
$\geq g(v)$
$=F_{d}(x,y^{*};u,v)$
$=f(y^{*})$
$\geq f(y)$
$=F_{d}(x,y;u,v),$
which completes the proof. ∎
The following lemma shows that extremal configurations collapse to dimension two.
Lemma 3.2 (Two-dimensional collapse).
Let $x,y,u,v\in\mathbb{R}_{>0}^{d}$ and set
$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
Assume that
$\frac{y_{i}}{x_{i}}\in\{a^{-},a^{+}\}\quad\text{and}\quad\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}\qquad\text{for all }1\leq i\leq d.$
Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm},$
and
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;\geq\;F_{d}(x,y;u,v).$
Proof.
If $a^{-}=a^{+}=a$, then $y=ax$ and $F_{d}(x,y;u,v)=1$. In this case one may take $x^{\prime}=(1,1)$, $y^{\prime}=(a,a)$, $u^{\prime}=(1,1)$ and $v^{\prime}=(b^{-},b^{+})$, so that $s^{\pm}(x^{\prime},y^{\prime})=a$, $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$, and $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=1$. The case $b^{-}=b^{+}$ is similar and is handled analogously. In either case the conclusion of the lemma holds.
We therefore assume from now on that $a^{-}<a^{+}$ and $b^{-}<b^{+}$.
Define the index sets
$I_{-}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{-}\,\},\qquad I_{+}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{+}\,\},$
which form a partition of $\{1,\dots,d\}$ and are both nonempty by assumption.
Step 1: Two–dimensional reduction preserving $F_{d}$.
Define
$x^{-}:=\sum_{i\in I_{-}}x_{i},\qquad x^{+}:=\sum_{i\in I_{+}}x_{i},\qquad x^{\prime}:=(x^{-},x^{+})\in\mathbb{R}_{>0}^{2},$
and
$u^{-}:=\frac{1}{x^{-}}\sum_{i\in I_{-}}x_{i}u_{i},\qquad u^{+}:=\frac{1}{x^{+}}\sum_{i\in I_{+}}x_{i}u_{i},\qquad u^{\prime}:=(u^{-},u^{+})\in\mathbb{R}_{>0}^{2}.$
Set
$y^{\prime}:=(a^{-}x^{-},\,a^{+}x^{+})\in\mathbb{R}_{>0}^{2}.$
By construction, $s^{\pm}(x^{\prime},y^{\prime})=a^{\pm}$. A direct computation shows that
$x\cdot u=x^{\prime}\cdot u^{\prime},\qquad y\cdot u=y^{\prime}\cdot u^{\prime}.$
Define an auxiliary vector $\tilde{v}\in\mathbb{R}_{>0}^{2}$ by
[…]
Then
$x\cdot v=x^{\prime}\cdot\tilde{v},\qquad y\cdot v=y^{\prime}\cdot\tilde{v}$,
and hence
[…]
Let
$\tilde{b}^{-}:=\frac{\tilde{v}^{-}}{u^{-}},\qquad\tilde{b}^{+}:=\frac{\tilde{v}^{+}}{u^{+}}.$
Then
$\tilde{b}^{-}=\frac{1}{x^{-}u^{-}}\sum_{i\in I_{-}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}},\qquad\tilde{b}^{+}=\frac{1}{x^{+}u^{+}}\sum_{i\in I_{+}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}}.$
Since $x^{-}u^{-}=\sum_{i\in I_{-}}x_{i}u_{i}$ and $\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}$, we conclude that $\tilde{b}^{-}$ is a weighted average of $b^{-}$ and $b^{+}$, and hence $\tilde{b}^{-}\in[b^{-},b^{+}]$. The same argument applies to $\tilde{b}^{+}$ and yields $\tilde{b}^{+}\in[b^{-},b^{+}]$.
Note that in general $\tilde{b}^{\pm}\neq b^{\pm}$ so we cannot simply take $v^{\prime}:=\tilde{v}$ and conclude the proof. An extra step is needed.
Step 2: Monotonicity argument.
For $b_{1},b_{2}>0$ define
$f(b_{1},b_{2}):=F_{2}\!\left(x^{\prime},y^{\prime};u^{\prime},(b_{1}u^{-},\,b_{2}u^{+})\right).$
Then
$f(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{Ab_{1}+Bb_{2}}{Cb_{1}+Db_{2}},$
where
$A:=a^{-}\,x^{-}u^{-},\qquad B:=a^{+}\,x^{+}u^{+},\qquad C:=x^{-}u^{-},\qquad D:=x^{+}u^{+}.$
Note that
$AD-BC=x^{-}u^{-}\,x^{+}u^{+}\,(a^{-}-a^{+})\;<\;0,$
since $a^{-}<a^{+}$. Then
$\frac{\partial f}{\partial b_{1}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(AD-BC)\,b_{2}}{(Cb_{1}+Db_{2})^{2}}\;<\;0,$
$\frac{\partial f}{\partial b_{2}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(BC-AD)\,b_{1}}{(Cb_{1}+Db_{2})^{2}}\;>\;0,$
for all $b_{1},b_{2}>0$. It follows that $f(b_{1},b_{2})$ is decreasing in $b_{1}$ and increasing in $b_{2}$.
Define
$v^{\prime}:=(b^{-}u^{-},\,b^{+}u^{+})\in\mathbb{R}_{>0}^{2},$
so that $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$. By construction,
$f(b^{-},b^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$
and
[…]
Since $\tilde{b}^{\pm}\in[b^{-},b^{+}]$ and $f$ is decreasing in its first argument and increasing in its second, we conclude that
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=f(b^{-},b^{+})\;\geq\;f(\tilde{b}^{-},\tilde{b}^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},\tilde{v})=F_{d}(x,y;u,v).$
This completes the proof.
Corollary 3.3 (Four-Point Collapse).
Let $d\geq 2$ and $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Then there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$s^{\pm}(x^{\prime},y^{\prime})=s^{\pm}(x,y),\qquad s^{\pm}(u^{\prime},v^{\prime})=s^{\pm}(u,v),$
and
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
Proof.
Apply Lemma 3.1 to obtain $y^{*},v^{*}$ with $F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y;u,v)$. Then apply Lemma 3.2 to $x,y^{*},u$ and $v^{*}$ and obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y^{*};u,v^{*})$. ∎
Theorem 3.4 (Envelope Inequality in Arbitrary Dimension).
Let $A\in\mathbb{R}^{d_{1}\times d_{2}}_{>0}$ and $B\in\mathbb{R}^{d_{2}\times d_{3}}_{>0}$, where $d_{1},d_{2},d_{3}\geq 2$. Then
$R(AB)\leq\Phi\bigl(R(A),R(B)\bigr).$
Proof.
By definition, $R(AB)$ is attained by a $2\times 2$ sub-block of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ attaining this maximum. Let
[…]
Then the corresponding $2\times 2$ block of $AB$ is
$\begin{pmatrix}x\cdot u&x\cdot v\\ y\cdot u&y\cdot v\end{pmatrix}$,
and therefore
$R(AB)=F_{d}(x,y;u,v).$
Set
$\alpha:=\operatorname{Dist}(x,y)\leq R(A),\qquad\beta:=\operatorname{Dist}(u,v)\leq R(B).$
By Corollary 3.3, there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
$\operatorname{Dist}(x^{\prime},y^{\prime})=\operatorname{Dist}(x,y)=\alpha,\qquad\operatorname{Dist}(u^{\prime},v^{\prime})=\operatorname{Dist}(u,v)=\beta,$
and
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then
$R(A^{\prime})=\operatorname{Dist}(x^{\prime},y^{\prime})=\alpha,\qquad R(B^{\prime})=\operatorname{Dist}(u^{\prime},v^{\prime})=\beta,$
and
$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix},\qquad\text{so}\qquad F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=R(A^{\prime}B^{\prime}).$
Using Theorem 2.2 and the monotonicity of $\Phi$,
$R(AB)$ $=F_{d}(x,y;u,v)$
$\leq F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$
$=R(A^{\prime}B^{\prime})$
$\leq\Phi\bigl(R(A^{\prime}),R(B^{\prime})\bigr)$
$=\Phi(\alpha,\beta)$
$\leq\Phi\bigl(R(A),R(B)\bigr).$
4 Distortion of Products
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and set
$P_{n}:=A_{n}\cdots A_{1},$
assuming the products are well defined.
It is convenient to work with the square root of the distortion,
$S(A):=\sqrt{R(A)}.$
The envelope inequality then takes the simpler form
$S(AB)\;\leq\;\Psi\bigl(S(A),S(B)\bigr),\qquad\Psi(p,q):=\frac{1+pq}{p+q}.$
Since $\Psi$ is increasing, distortion bounds propagate along products. If
$R(A_{n})\leq\alpha_{n},\qquad\text{i.e.}\qquad S(A_{n})\leq p_{n}:=\sqrt{\alpha_{n}},$
then
$S(P_{n})\leq q_{n},$
where $(q_{n})$ is defined recursively by
$q_{1}=p_{1},\qquad q_{n+1}=\Psi(p_{n+1},q_{n}).$
Finally,
$R(P_{n})\leq q_{n}^{2}.$
Uniformly bounded distortions.
Assume that all matrices satisfy the uniform bound
$R(A_{n})\leq\alpha,\qquad\text{i.e.}\qquad S(A_{n})\leq p:=\sqrt{\alpha}.$
Then the upper bounds $q_{n}$ satisfy
$q_{n+1}=\Psi(p,q_{n}),\qquad q_{1}=p.$
This recursion is the iteration of a single Möbius map. Writing
$\kappa:=\frac{p-1}{p+1},$
we obtain the closed-form expression
$q_{n}=1+\frac{2\,\kappa^{n}}{1-\kappa^{n}}.$
Consequently,
$R(P_{n})\leq q_{n}^{2}=1+\frac{4\,\kappa^{n}}{(1-\kappa^{n})^{2}}.$
Collecting the preceding arguments, we record the following theorem.
Theorem 4.1 (Finite-Product Distortion Bound).
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices, possibly of varying dimensions, and assume that all products $P_{n}:=A_{n}\cdots A_{1}$ are well defined.
For any positive matrix $B>0$, define its distortion by
$R(B):=\max_{i,j,k,\ell}\frac{b_{ik}\,b_{j\ell}}{b_{i\ell}\,b_{jk}}.$
Assume that all factors satisfy the uniform bound
$R(A_{n})\leq\alpha\qquad\text{for all }n\geq 1.$
Set
$\kappa:=\frac{\sqrt{\alpha}-1}{\sqrt{\alpha}+1}.$
Then, for every $n\geq 1$,
$R(P_{n})\;\leq\;1\;+\;\frac{4\,\kappa^{\,n}}{(1-\kappa^{\,n})^{2}},$
and therefore
$R(P_{n})\longrightarrow 1\quad\text{at least exponentially fast as }n\to\infty.$
5 Comparison with the Birkhoff–Bushell
Contraction Theory
In classical Birkhoff–Bushell (BB) theory one studies distortion on a log scale via the Hilbert metric
$d_{H}(x,y):=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
If $B$ is an $n\times 2$ matrix with columns $x$ and $y$, and $A$ is any positive $n\times n$ matrix, then
$d_{H}(x,y)=\log R(B),\qquad d_{H}(Ax,Ay)=\log R(AB).$
In this logarithmic setting, the envelope inequality takes the form
$d_{H}(Ax,Ay)\;\leq\;\Theta\!\left(d_{H}(x,y)\right),$
where
$\Theta(h):=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p:=\sqrt{R(A)}.$
A direct computation gives
$\Theta^{\prime}(h)=\frac{(p^{2}-1)e^{h/2}}{(1+pe^{h/2})(p+e^{h/2})},\qquad\Theta^{\prime\prime}(h)=-\frac{(p^{2}-1)e^{h/2}p(e^{h}-1)}{2(1+pe^{h/2})^{2}(p+e^{h/2})^{2}}.$
In the nontrivial case $p>1$, it follows that $\Theta$ is increasing and concave on $0,\infty)$, and satisfies
$\Theta(0)=0,\qquad\Theta^{\prime}(0)=\frac{p-1}{p+1}=:\kappa(A).$
By concavity, $\Theta$ lies below its tangent line at the origin,
$\Theta(h)\;\leq\;\kappa(A)\,h,$
which yields the classical BB contraction estimate
$d_{H}(Ax,Ay)\;\leq\;\kappa(A)\,d_{H}(x,y).$
Moreover,
$\lim_{h\to\infty}\Theta(h)=\log R(A),$
so the envelope bound remains uniformly bounded, whereas the linear BB bound $\kappa(A)h$ diverges. The two bounds therefore agree to first order at $h=0$, but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp for all $h$.
Figure 2 compares the nonlinear envelope $h\mapsto\Theta(h)$ with the classical BB line $\kappa(A)h$ and the saturation level $\log R(A)$.
[…]
Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$. The envelope saturates at $\log R(A)$, while the BB bound grows without bound.